# Cocoa: Co-Planning and Co-Execution with AI Agents

### v2

@@ 5.5. Data Analysis
The first author analyzed the recording transcripts using Braun and Clarke’s reflexive thematic analysis (Braun and Clarke, 2019). This approach uses a hybrid inductive-deductive approach to iteratively surface codes and themes across the data. We paid close attention to codes related to participants’ comparisons of their experiences across the two systems and gradually grouped them into themes. NotebookLM [^10] https://notebooklm.google.com was used to help iterate on themes and discover new ones. We performed statistical tests of participants’ ratings on the evaluation forms with the Wilcoxon signed-rank test (with the Bonferroni correction for our multi-rating analysis on significant results) given the non-parametric nature of our data. We also saved all documents used in the study as well as all conversation histories in the baseline system, and referred back to them for context when needed.
Figure 12. An overview of participants’ Likert scale ratings of steerability, ease of use, and general utility across our baseline and Cocoa.
6. Results
@@
In this section, we present quantitative and qualitative analyses of participants’ data from our user study. We group these results by our research questions from the start of Section 5.
6.1. Steerability, Ease of Use, and Utility (RQ1)
@@
Introducing novel interactive systems with additional affordances can often lead to higher effort when using the systems (Shneiderman and Maes, 1997; Hutchins et al., 1985). However, based on participants’ post-task ratings on a 5-point Likert scale, there was no significant difference in perceived ease of use between the baseline and Cocoa with N participants ($M_{b}=4.5,M_{c}=4,p=1.000,W=7.50$). The distributions of ratings between the two conditions were nearly identical, as indicated by $p=1.000$. At the same time, we observed a significant difference that Cocoa provided better steerability. In response to the question “I could easily steer the system towards doing something helpful,” participants rated Cocoa higher than the baseline to a significant degree ($M_{b}=3,M_{c}=4,p=0.005,W=0$.) [^11] Here, we set the significance threshold to $\alpha=0.05/3=0.015$ using Bonferroni correction. Our result is significant post-correction: $p=0.005<0.015$ $W=0$ indicates that all participants rated Cocoa’s steerability as greater than or equal to that of our baseline. In sum, these results suggest that Cocoa provided rich and flexible affordances that can improve steerability of AI agents without sacrificing ease of use (12).
We also tested for differences in general utility, which was a composite metric consisting of three measures broadly related to the usefulness and insightfulness of system outputs, but we found no significant differences ($M_{b}=M_{c}=3.5,p=0.7,W=40$). This was unsurprising—we did not explicitly aim to improve general utility (but hoped it would not decrease) as the systems’ perceived utility may depend on a range of system-agnostic factors beyond our control (e.g., contextual interpretations of outputs by researchers, researchers’ existing understanding of the problem space, etc.) However, improved steerability may have allowed users to steer the agent away from providing downright unhelpful outputs (as indicated by the fewer ratings of 1 and 2 for Cocoa in Fig. 13). Additionally, the nature of the task may have influenced perceived utility, as Cocoa might be better suited for certain tasks, while chat may work better for others. The random assignment of tasks in our study procedure can thus limit how informative quantitative measures of utility can be when evaluating the two systems. Qualitative insights from post-task interviews and participants’ think-aloud sessions can provide deeper insights into the task suitability of interactive plans, and we unpack this in Section 6.2.
Figure 13. A breakdown of participants’ Likert scale ratings of steerability, ease of use, and general utility across our baseline and Cocoa. The three measures included in general utility are highlighted in blue.
@@ 6.1.1. Co-planning afforded steerability
Participants found various co-planning features provided by Cocoa helped improve steerability over and control of the agent. Many thought that the ability to compose, edit, and rerun plan steps allowed them to better “fine-tune” the agent’s outputs and their research process. P1 said that they could “use [plan steps] as building blocks” to create a custom workflow. In addition, they wished that even more plans were presented to them in the plan selector and to interactively drag and drop steps across multiple plan versions to compose them.
P7 and P15 both appreciated the ability of edit and rerun steps, as it enabled them to quickly “organize thoughts and iterate on ideas” (P7); P15, specifically, iteratively adjusted a “paper search” step, reran it to cover papers from venues that they preferred, and was able to obtain more relevant final outputs with running the rest of the steps in the plan. Another interesting strategy was that P16 found “backtracking” to a specific step and editing it to be much more helpful than digging through a chat conversation: “With chat, when does something wrong there’s not really an easy way to fix it because there’s no concrete steps that it’s following, whereas [with Cocoa ], I can go back and be like, let me have it do something else here.” Another interesting use case was that P12 added a step at the end of their plan to have the agent suggest some potential next steps to pursue after seeing its paper summaries. [^12] Cocoa does not yet support nested planning; otherwise P12 could have generated an interactive plan based on the last plan step by invoking the agent on it.
Besides iterating and steps by editing and rerunning them, we also observe participants removing system-suggested steps to improve the plans. For example, P5 removed a less relevant intermediate step on searching for papers related to autonomous vehicles to avoid distractions from this tangential topic. In sum, participants leverage different features to co-plan with Cocoa, including adding, removing, editing, re-running, and composing steps in the plan to explore different execution strategies while staying focused to the original goals they set out to achieve.
@@ 6.1.2. Co-execution afforded steerability
Participants liked having the flexibility to choose between stepwise and continuous plan execution. Many chose stepwise execution because it allowed them to better control the overall process and prevent error propagation. P4 and P11 both wanted to verify and manually curate the list of papers returned by the agent from a paper search step before the agent moves onto subsequent steps, with P11 specific citing a resource efficiency as a practical reason: “If I update [outputs of] the first or second step it would need to rerun the following steps. I just want to reduce some API usage.” P10 similarly mentioned that they would “waste energy” running subsequent steps if they detected errors in initial steps. P3 and P10, who were both computational notebook researchers, both made connections between stepwise execution to notebooks. P3 said that the workflow was familiar and aligned with theirs: “I wanted more like a Jupyter notebook, step by step [process]. I think that that was just very salient to how I think.” P10 said that their choice was partly due to “residual habits” from using computational notebooks and wanting to evaluate each step’s output before proceeding.
However, a few participants preferred continuous execution. P8 reasoned that continuous execution was more efficient because it parallelized human and AI efforts: “Running all is a little bit of parallelization in my head where I can get the next thing rolling, and if I end up not having to edit it, that means we’re already moving ahead.” P9 echoed this sentiment: “I can then look at the intermediate results and then wait for the others to keep running, since it might take some time for the system to run each step.” P15 stated that while they opted for continuous execution, their choice depended on the familiarity with the task: “if it were a topic that I felt like I didn’t already know what type of literature I would be looking for or what the output should look like, I would go through them individually and maybe use it as more of a literature exploration tool. But in this case, I already know in general what types of papers I want to see it come up with.” This underscores the importance of providing users with a choice of execution method to flexibly cater to researchers’ context-specific needs and preferences.
@@
6.1.3. Cocoa was perceived as more transparent than chat.
Participants shared that Cocoa was more transparent due to its clear display of step and plan completion statuses and information provenance, which may have contributed to Cocoa’s steerability. P3 enjoyed being able to control and run each step in Cocoa: “I did like the step by step because that gave me here’s what’s running when and in what order, versus the chat is going to be more of a blackbox.” They also greatly valued provenance and liked how the agent left a trace of “what process is running.” Supporting co-execution of individual steps also positively impacted transparency—when conducting literature surveys, P12 preferred Cocoa because it allowed them to “see which papers were being used to generate content, and edit those papers as I see fit, which isn’t really directly possible to do in chat.” For both P6 and P7, the visual step-by-step format of Cocoa provided a valuable, interactive trace of their past workflows with the agent, which also provided a clearer view of the agent’s actions. P11 said that having clear indicators of where the agent was in the plan and surfacing that information when the plan is collapsed was “really convenient when you have several agents running at the same time.”
6.2. Task-Specific Preferences for Interactive Planning (RQ2)
It is reasonable (and expected) for Cocoa to not outperform our chat baseline for every type of research task. After all, chat is a strong baseline because it is flexible, easy to use, and ubiquitous. Instead, in this section, we aim to surface participants’ insights on when interactive plans may be preferred over chat.
Participants generally agreed that when tackling questions or tasks that required a structured or organized approach, they preferred Cocoa over chat. One example of this type of task is literature review and synthesis. Among other participants, P12 considered the ability to transparently see and edit papers (i.e., in the output of a plan step) throughout a structured plan especially important when conducting literature surveys. They also considered the plan more usable as an end product than a conversation: “I think it’s a better way to have a finished draft [plan] that I can directly use as opposed to trying to dilute a plan in a conversation.” P7 felt more organized with Cocoa in information-intensive tasks such as literature review because the agent outputs, their notes, and plans are all in one place with provenance: “everything can be part of one consolidated document which I can refer to again.” P14 agreed, saying the plan also acted as a “really good way to organize my thoughts.” P1 found it helpful that in the initial steps, Cocoa could cast a wide net for papers using diverse search queries and allowed them to do a ‘‘rough filtering of papers” that may be relevant but outside of their immediate area. In general, the plans created by the agent for literature review aligned with researchers’ typical workflows: P15 likened the plan to be “what I would recommend to a junior student.” Because of this alignment, some participants such as P9 and P10 found it efficient for the agent to take over the “really banal parts of having to [brainstorm] and type out a plan” (P10).
The structure of an interactive plan was also helpful when participants may have an idea of their desired final output, but were unsure of how to get there. For their task, P6 appreciated Cocoa because they needed “some concrete steps even if I need to iterate over those steps,” and that the agent’s first step “helped me get to search terms [for papers] which was what I was struggling with.” P14 shared this sentiment, saying that plans can help them overcome mental blockers when faced with a tough problem: “If I’m really stuck frustrated, I can click a single button it could output a whole plan for me. That would make me feel really safe because it’s always here outputting something for me which potentially get me out of this frustration.” For P16, they did not consider anything but a structured approach to be desirable for research. They envisioned the research process as “a lot of subtasks that you just naturally kind of come up with” and knowledge they collected to be “concrete items” on which they could “do high-level actions” such as extraction. As a result, Cocoa aligned well with their mental model and found the chat “very frustrating […] and almost like the opposite” of their preferred approach. Some others, such as P3 and P10, simply did not want to engage in chat-based interactions during the research process due to hallucinations in text-only responses (P3) and finding the conversational workflow to be unnecessary and inefficient in research (P10, P16).
While interactive plans afford more steerability for users when interacting with AI agents, we also observe that our chat baseline interface may be more ideal for particular categories of tasks. One category of tasks we observed involved freeform exploration, where the nature of the desired final output was ambiguous. For example, P2, while conducting an early brainstorming task, thought it was “a little overwhelming to go through all these steps and I also felt this kind of pressure to stick with the plan although I was editing it.” P1 further explained when brainstorming, “any kind of structure is a barrier to thinking deeply” and that the plan, while helping them stay on track in finishing their task, made them feel constrained. Instead, they thought that Cocoa should behave closer to a chat interface to align better with their brainstorming workflow, proposing that Cocoa could show them “a single step at a time” and allowed them to “choose from options for the next step” after they have seen the outputs for the previous step. Another category of tasks pointed to by the participants was narrowly scoped question-answering—specifically, P13 saw chat as better UIs for “question answering [system]” where the main interaction modality is to “keep on asking [new and follow-up] questions.”
@@ 6.2. Task-Specific Preferences for Interactive Planning (RQ2)
At a high level, a chat UI may be more appropriate when the task at hand is open-ended and difficult to create plans in advance, or too narrowly scoped to benefit from a multi-step plan. In the end, many participants recognized that both interfaces have their strengths and weaknesses, and, in P13’s words, it was “not fair to say one is better than the other.” Instead, participants imagined using the two systems at different stages of their research. P5 compared their experiences between the two systems as follows: “in order to make [Cocoa ] do the types of tasks that I want it to do, I need to have a more specific idea of what I want it to do.” In Section 7.1, we further discuss future design implications, including the costs and benefits between the two paradigms and potential ways of combining them to get the best of both worlds.
6.3. Task Assignment to Agent vs. User (RQ3)
In our formative study (Section 3.3.3 and 3.3.4), participants shared that they preferred to assign tasks related to information retrieval and brainstorming tasks to the agent while keeping higher-level reasoning and synthesis tasks for themselves. In practice, however, most participants toggle all plan steps—even higher-level ones assigned to them by default—to the agent. Those who left some steps as user steps toggled the step to the agent when it arrived at that step. In this section, we use our qualitative data to discuss two reasons for this—assessing agent capability and time constraints—and how participants imagined the assignment to be different with prolonged use.
6.3.1. Participants mostly used agent steps during the study
@@
Some participants did not choose to include user steps in their plans because they were curious about the agent’s capabilities and limitations—a natural tendency when learning to use a new system. For example, P1 wanted to see “how the bot does it” while acknowledging that they prefer to “make the decisions and here, these are more exploratory steps rather than decision-making ones.” P13 also opted for this approach and considered the agent’s errors to be low-stakes because they can directly modify the outputs at any time: “when I don’t like something, I can just easily remove it.” P9 also felt assured that they could edit the outputs and did not consider the output quality to be particularly problematic: “usually as long as it’s somewhat helpful and it’s just running in the background, I would take a look at [the output].” Adopting a cost-benefit framework, they mentioned that they would leave a step as an agent step as long as “the cost of looking at the agent’s output is less than the benefit of whatever insights I obtain.” Participants were also influenced by the time constraints of the study. P8 explicitly said that “I assign everything to [the agent] because I don’t think we have that much time for me to refine my [outputs].” Some participants were not as explicit about this as P8, but the study’s environment and structure—combined with the system’s output editing affordances—likely encouraged many participants to direct the agent to make an initial attempt at completing the plan.
6.3.2. Step assignment could change over longer-term use
Many participants admitted that their step assignments may change if they were to use Cocoa for an extended period of time outside of the study environment. In particular, they imagined user steps to be tasks involving higher-level synthesis and creative conceptualization. These tasks often had intrinsic value to participants that would be given up if they were automated. For example, P2 said they would assign synthesis tasks to themselves because it “can help me specify or narrow down the context to what I’m interested into myself.” Going further, P15 discussed the importance of “staying close to the data” throughout the research process, particularly for more junior researchers:
“I really believe that science happens through the interaction between the researcher and the data at hand. And I think it’s a really important part of just scholarship in general, especially for junior scholars, to do that synthesis on their own. That’s how they learn how to do science and how to make a contribution.” [P15]
They also questioned the ability of AI to make meaning out of research data: “At the end of the day, [AI]’s doing nothing more than just analyzing what topics and words and phrase patterns are in the paper data set that it has. It’s not like interacting with the data in the same way that a researcher interacts with the data.” P10, who shared that they are “the kind of person to just follow the [agent’s] plan,” pointed out that that excessive automation can dull their research intuition by removing the need to read papers themselves: “Maybe I would not have gone through and read all of the papers. From past experience that’s sometimes not going to be the most effective way to work.” For P13, the desire to maintain some agency to guide the agent is a great reason to add in user steps before the agent searches for literature: “a very good use case for me would be to give [the agent] 3 papers and have it tell me, ok, these are newer papers related to it.” Similarly, if given more time, P5 planned to provide some starter papers in their area upon seeing that the agent misinterpreted an earlier search attempt: “I should give it more specific papers. I’m worried it’ll get a little confused by all these [other papers].” Longer-term use of Cocoa may surface step assignment preferences more similar to those from our formative study.
7. Discussion and Future Work

### v3

@@ 5.4. Data Analysis
The first author analyzed the recording transcripts using Braun and Clarke’s reflexive thematic analysis (Braun and Clarke, 2019). This approach uses a hybrid inductive-deductive approach to iteratively surface codes and themes across the data. We paid close attention to codes related to participants’ comparisons of their experiences across the two systems and gradually grouped them into themes. NotebookLM [^8] https://notebooklm.google.com was used to help iterate on themes and discover new ones. We performed statistical tests of participants’ ratings on the evaluation forms with the Wilcoxon signed-rank test (with the Bonferroni correction for our multi-rating analysis on significant results) given the non-parametric nature of our data. We saved all documents used in the study as well as all conversation histories in the baseline system, and referred back to them when needed.
The first author also coded all video recordings from the study to note the timestamps at which participants engaged and disengaged in a particular interaction. In Cocoa, these interactions were co-planning (see 4.1), co-execution (see 4.2), and output inspection (when the participant passively inspected the system’s outputs). The closest equivalent interactions were used as codes in the baseline: prompting for planning (instructing the agent to perform an action without any attempts to modify the agent’s earlier outputs), prompting for editing [^9] Since there were no user steps in the baseline, the closest equivalent interaction is attempting to edit the agent’s output via prompting. (instructing the agent to modify or only consider a subset of its output in future actions), and output inspection. During video coding, the first author also captured the number of input items (items generated as part of the output of a plan step or conversation turn) and output items (items fed as context into the next plan step or conversation turn) for all plan steps or turns that output discrete items (i.e., papers, authors, topics, text entities).
6. Lab Study Results
A high-level overview of participants’ interactions with Cocoa and the baseline is depicted in Figure 12. In both conditions, output inspection occupied the most time out of the coded interactions, but was lengthier in the baseline than Cocoa (54.1% vs. 33.6% of the approximately 25-minute session). A paired t-test found this difference to be significant ($t(15)=3.95,p<0.001$). Participants also engaged in co-execution in Cocoa significantly more often than the closest equivalent interaction—output editing via prompting—in the baseline (15.2% vs. 3.1%, $t(15)=4.29,p<0.001$). They spent slightly more time on co-planning (21.5% vs. 17.4%), although this was not significant ($t(15)=1.01,p=0.33$). Our analysis indicates that, when provided with more affordances to interactively modify the agent’s outputs, users will take advantage of those opportunities for active engagement and shift away from passive output consumption.
Figure 12. Participants’ interaction activity in Cocoa and the baseline within the study sessions. *** indicates a statistically significant difference ($p<0.001$ via a paired t-test) with the closest equivalent interaction in the opposing condition. Gray areas represent spans of time where the participant was not engaging in any of our coded interactions because they were answering a question posed by the study facilitator and/or waiting on the system. Definitions for all coded interactions can be found in Section 5.4.
[…]
A visualization of participants’ interaction activity in Cocoa (left) and the baseline (right). Each participant on the y-axis is represented as a bar that is divided into chunks of co-planning, co-execution, and output inspection for Cocoa, and prompting (planning), prompting (editing) and output inspection for the baseline. The x-axis depicts normalized session duration from 0.0 to 1.0.
Could more output editing activity mean that participants are less satisfied with Cocoa’s outputs? To answer this, we analyzed the output-to-input item ratio for each round of interaction—where a round is a plan step in Cocoa or conversation turn in the baseline—across all participants. The output-to-input ratio tells us what percentage of items generated by a round are carried over as context to be used in the next round. As can be seen in Figure 13, the ratio is much higher in Cocoa ($M=97.5,\sigma=23.0$) compared to the baseline ($M=0,\sigma=33.0$). This difference is significant per a Mann-Whitney U Test ($U=1722.0,p<0.001$). These results suggest that the flow of data within the two systems differ fundamentally: the flow of items from one plan step to the next is rather fluid in Cocoa where much results were carried over to the next steps in the plan, but disjoint in the baseline where most or all of the results were unused. This suggests that participants, when using the baseline system, more often focused on few items in the output, or pivoted to a different task entirely. Thus, there is no indication that increased output editing activity is due to dissatisfaction with system outputs.
Figure 13. The output-to-input ratio in Cocoa and the baseline, across all rounds of interaction from all participants ($n=16$). *** indicates $p<0.001$ via a Mann-Whitney U Test.
[…]
A boxplot depicting the output-to-input ratio in Cocoa and the baseline. Cocoa is shown above the baseline on the y-axis and has a median of 97.5, whereas the baseline has a median of 0. The output-to-input item ratio is shown on the x-axis from 0 to 100. The difference between the two conditions is highly significant.
@@ 6. Lab Study Results
In the rest of this section, we present quantitative and qualitative analyses of participants’ data from our user study. We group these results by our research questions from the start of Section 5.
Figure 14. Participants’ Likert scale ratings of steerability, ease of use, and general utility across our baseline and Cocoa. ** indicates $p<0.01$ via a Wilcoxon signed-rank test.
[…]
Three box plots showing participants’ Likert scale ratings of steerability, ease of use, and general utility across Cocoa and the baseline. Cocoa is shown above the baseline on the y-axis in all 3 plots. All 3 plots have an x-axis depicting the Likert scale rating from 1 to 5. Participants rated Cocoa’s steerability as significantly higher compared to the baseline.
6.1. Steerability, Ease of Use, and Utility (L1)
@@
Introducing novel interactive systems with additional affordances can often lead to higher effort when using the systems as a trade-off for better utility (Shneiderman and Maes, 1997; Hutchins et al., 1985). However, based on participants’ post-task ratings on a 5-point Likert scale, there was no significant difference in perceived ease of use between the baseline and Cocoa with 16 participants ($M_{b}=4.5,M_{c}=4,p=1.000,W=7.50$). At the same time, we observed a significant difference that Cocoa provided better steerability. In response to the question “I could easily steer the system towards doing something helpful,” participants rated Cocoa higher than the baseline to a significant degree ($M_{b}=3,M_{c}=4,p=0.005,W=0$), [^10] Here, we set the significance threshold to $\alpha=0.05/3=0.015$ using Bonferroni correction. Our result is significant post-correction: $p=0.005<0.015$. with $W=0$ indicating that all participants rated Cocoa’s steerability as greater than or equal to that of our baseline. In sum, these results suggest that by taking inspiration from familiar interaction patterns of computation notebooks and text editors, Cocoa provided richer affordances, and does so without sacrificing ease of use to better support human-AI collaboration compared to the simple chat interface used in the baseline condition (Figure 14).
We also tested for differences in general utility, which was a composite metric consisting of three measures broadly related to the usefulness and insightfulness of system outputs, but we found no significant differences ($M_{b}=M_{c}=3.5,p=0.7,W=40$). This was unsurprising— the systems’ utility to the researcher may depend on a range of system-agnostic factors beyond our control (e.g., contextual interpretations of outputs by researchers, the extent to which researchers thought about the task before the study, etc.) However, improved steerability may have allowed users to steer the agent away from providing downright unhelpful outputs in Cocoa, as indicated by fewer lowly-rated outliers (see Figure 14).
Additionally, the nature of the task may have influenced perceived utility, as Cocoa might be better suited for certain tasks, while chat may work better for others. Qualitative insights from post-task interviews and participants’ think-aloud sessions provide insights into the task suitability of interactive plans (see Section 6.2.1).
@@ 6.1.1. Co-planning afforded steerability
Qualitative results based on think alouds and post-interviews showed that participants found various co-planning features provided by Cocoa helped improve the steerability of the agent. Many thought that the ability to compose, edit, and rerun plan steps allowed them to better “fine-tune” the agent’s outputs and their own research process. P1 was inspired to “use [plan steps] as building blocks” to create a custom workflow. P7 and P15 both appreciated the ability to edit and rerun steps, which enabled them to quickly “organize thoughts and iterate on ideas” (P7); P15, specifically, iterated on “paper search” step, reran it to cover papers from venues that they preferred, and was able to obtain more relevant final outputs with running the rest of the steps in the plan. Another strategy was “backtracking” to a specific step and editing it which P6 found to be much more helpful than digging through a chat conversation: “With chat, when it does something wrong there’s not really an easy way to fix it because there’s no concrete steps that it’s following, whereas [with Cocoa], I can go back and be like, let me have it do something else here.”
Besides iterating and steps by editing and rerunning them, we also observed participants adding and removing agent-suggested steps to improve the plans. For example, P12 added a step at the end of their plan to have the agent suggest some potential next steps to pursue after seeing its paper summaries, while P5 removed a less relevant intermediate step on searching for papers related to autonomous vehicles to avoid distractions from this tangential topic.
@@ 6.1.2. Co-execution afforded steerability
Participants appreciated the flexibility to choose between stepwise and continuous plan execution. Many chose stepwise execution because it allowed them to better control the overall process and prevent error propagation. P4 and P11 both wanted to verify and manually curate the list of papers returned by the agent from a paper search step before proceeding to subsequent steps. P11 specifically cites resource efficiency as a practical reason: “If I update [outputs of] the first or second step it would need to rerun the following steps. I just want to reduce some API usage.” P10 similarly mentioned that they would “waste energy” running subsequent steps if they detected errors in initial steps. Others preferred continuous execution. P8 and P9 thought continuous execution was more efficient because it parallelized human and AI efforts: “Running all is a little bit of parallelization in my head where I can get the next thing rolling, and if I end up not having to edit it, that means we’re already moving ahead” (P8). P15 stated that while they opted for continuous execution, their choice depended on the familiarity with the task: “if it were a topic that I felt like I didn’t already know what type of literature I would be looking for or what the output should look like, I would go through them individually and maybe use it as more of a literature exploration tool. But in this case, I already know in general what types of papers I want to see it come up with.” This underscores the importance of providing users with a choice of execution method to cater to researchers’ contextual needs and preferences.
@@
6.2. Task-Specific Preferences for Interactive Planning vs. Chat (L2)
@@
It is reasonable for participants not to strictly prefer Cocoa over our chat baseline for every research task. After all, chat is a strong baseline because it is flexible, easy to use, and ubiquitous. Here, we aim to surface participants’ insights on when they prefer interactive plans over chat, and vice versa. Ultimately, many participants recognized that both interfaces have their strengths and weaknesses, and, in P13’s words, it was “not fair to say one is better than the other.” Instead, participants imagined using the two systems at different stages of their research. In Section 8.1, we further discuss the costs and benefits between the two paradigms and how to combine them in future systems to get the best of both worlds.
6.2.1. Preferences for interactive plans
Participants generally agreed that when tackling questions or tasks that required a structured or organized approach, they preferred Cocoa over chat. One example of such tasks is literature review and synthesis. P12 considered the plan more usable as an end product than a conversation: “I think it’s a better way to have a finished draft [plan] that I can directly use as opposed to trying to dilute a plan in a conversation.” Others appreciated that the plan helped them stay organized and consolidated lot of information: “[the plan] is a really good way to organize my thoughts” (P14) and “everything can be part of one consolidated [artifact] which I can refer to again” (P7).
The structure of an interactive plan was also helpful when participants had an idea of their desired final output, but were unsure of how to get there. For their task, P6 appreciated Cocoa because they needed “some concrete steps even if I need to iterate over those steps,” and because the agent’s first step “helped me get to search terms [for papers] which was what I was struggling with.” Similarly, P14 said that plans can help them overcome mental blockers when faced with a tough problem: “If I’m really stuck and frustrated, I can click a single button, and it could output a whole plan for me. That would make me feel really safe because it’s always here outputting something for me which potentially get me out of this frustration.” For P16, they did not consider anything but a structured approach to be acceptable for research. They envisioned the research process as “a lot of subtasks that you just come up with” and collected information to be “concrete items” that they would operate on with “high-level actions [like] directly extracting insights.” As a result, Cocoa aligned well with their mental model and found the chat “very frustrating […] and almost like the opposite” of their preferred approach.
6.2.2. Preferences for chat-based interactions
While interactive plans afford more steerability, chat-based interactions may be more appropriate when the task at hand is open-ended and difficult to create plans in advance, or too narrowly scoped to benefit from a multi-step plan. Many preferred chat for freeform exploration, where the nature of the desired final output was ambiguous. P2, while conducting an early brainstorming task, thought it was “a little overwhelming to go through all these steps, and I also felt this kind of pressure to stick with the plan although I was editing it.” P1 suggested that, when brainstorming, Cocoa could show them “a single step at a time” and allow them to “choose from options for the next step” after they have seen the outputs for the previous step. Question-answering was another category of tasks where chat was preferred, where the next step depends on the output of the previous step. P13 found it much easier to “keep on asking [follow-up] questions” to drill down on a specific line of inquiry in chat.
6.3. Task Assignment to Agent vs. User (L3)
@@ 6.2.2. Preferences for chat-based interactions
In our formative study (Section 3.2.3), participants shared that they preferred to assign tasks related to information retrieval and brainstorming tasks to the agent while keeping higher-level reasoning and synthesis tasks for themselves. In our study, however, most participants toggle all plan steps—even higher-level ones assigned to them by default—to the agent. Some participants chose not to include user steps in their plans because they were curious about the agent’s capabilities and limitations—a natural tendency when learning to use a new system. For example, P1 wanted to see “how the bot does it” while acknowledging that they prefer to “make the decisions and here, these are more exploratory steps rather than decision-making ones.” P13 also opted for this approach and considered the agent’s errors to be low-stakes because they can directly modify the outputs at any time: “when I don’t like something, I can just easily remove it.” P9 also felt assured that they could edit the outputs and did not consider the output quality to be particularly problematic: “usually as long as it’s somewhat helpful and it’s just running in the background, I would take a look at [the output].” Adopting a cost-benefit framework, they mentioned that they would leave a step as an agent step as long as “the cost of looking at the agent’s output is less than the benefit of whatever insights I obtain.”
While we did not seen substantial use of user steps in the lab, participants also did not view the ability to create user steps negatively. Many speculated in the interviews that they would leverage this mechanism more if they had more time to interact with the system and outside of the lab environment. P8 explicitly said that “I assign everything to [the agent] because I don’t think we have that much time for me to refine my [own steps].” This further motivated us to conduct a longitudinal deployment study of Cocoa, and Section 7.2.1 details how participants’ task assignment strategy differs in the field.
7. Deployment Study and Results
As surfaced in Section 6.3, participants’ usage behaviors in Cocoa may shift given more time to interact with the system. To obtain a more holistic understanding of how Cocoa can support researchers in their day-to-day work, we conducted a 7-day field deployment study with 7 participants.
7.1. Participants and Procedure
We invited all participants from our lab study to participate in this field deployment. 7 participants (P3, P5, P10, P11, P12, P15, P16) agreed to participate.
Participants used Cocoa to work on a real-world, in-progress research project of their choice over a 7-day period. Before the study, the study facilitator sent each participant a user manual documenting Cocoa’s functionalities. The manual also contained some general guidelines for the week (see Appendix J), including aiming to spend at least 90 minutes with the system throughout the duration of the study. Each participant could log onto their personal workspace in Cocoa with a set of credentials provided by the study facilitator. We used system logs to verify that there was at least semi-regular activity for each participant throughout the study.
At the end of the 7-day period, participants completed a 30-minute semi-structured exit interview where they were asked about their general experience with the system as well as specific co-planning and co-execution interactions the study facilitator identified from within their workspace. All interviews were recorded, transcribed, and analyzed with the same qualitative analysis procedure described in Section 5.4. Participants received a \$100 USD honorarium upon completing the exit interview.
7.2. Field Deployment Study Results
7.2.1. Step assignment strategy changed: more user steps for complex longitudinal tasks
Compared to the lab study, where most co-execution was focused on removing or adding to an agent step’s output, participants in the field study assigned more plan steps to themselves. These user steps often involve high-level strategic decision-making, writing and developing arguments, running experiments and evaluations, deep reading of literature, and articulating core research ideas and novelty. The ability to assign steps to the agent and themselves helped participants create a back-and-forth workflow with the agent. For example, P15 first figured out “broad buckets of literature or themes or topics that I need to talk about” before passing those to the agent to “help me brainstorm ways to connect those bodies of literature.” P3, who had already conducted some interviews for their project, first had the agent look up key literature before assigning themselves a step to “pass in more of the dynamics from my interview notes” and then had the agent propose interventions based on the literature and notes. For many participants, user steps came later in the plan because of their high-level nature. P11 often assigned the last plan step to themselves because “it was about the main innovation in the project idea […] no one has [done it] before so I don’t trust the agent.” P16 assigned themselves the last step as a to-do item and a reminder to reason about how the agent’s output “could actually be tied into some of the document.” Participants also indicated that their step assignment also depended on the stage of research they were in. P3 admitted that “if I was just doing initial discovery and study planning, I would have gone more agent [steps].”
7.2.2. Interleaving co-planning and co-execution was valuable
Participants identified the interleaving of co-planning and co-execution afforded by interactive plans as a unique and valuable aspect of Cocoa. This interleaving allowed for a flexible and iterative workflow. P12 shared that “I don’t need to run the [earlier] retrieval step again. I know that those papers are relevant” but instead could focus on “only improving the second step […] to identify something different from the literature.” Similarly, P16 felt like they could easily “run the plan fully and then go in and adjust the rerun things,” which enabled them to “debug” suboptimal parts of the plan: “by forcing these checkpoints of brainstorming search queries and assembling lists of papers, it helps me get a sense of what parts of the process are good and which parts need work.” Outputs from co-execution also inspired new approaches for participants to “iteratively build upon a plan or reshape a plan” (P5). Participants even mentioned that they wanted similar interleaving strategies integrated into other LLM-powered systems; for P11, it was “one of the improvements I want to see in LLM [reasoning] models” so that users can force an early stop to the reasoning or steer it in more productive directions.
7.2.3. Cocoa was most useful for literature synthesis and early-stage project planning
All participants agreed that Cocoa was especially helpful for literature discovery and synthesis. P15 found the plan steps involving literature search “super useful” and the search query generation steps that commonly preceded them to be “really helpful for refining my thought process because it made me think more carefully about what I’m actually trying to look for.” P10 also found Cocoa to be “really helpful with looking for papers” and that it was “less burdensome” to use compared to Google Scholar. The combination of plan steps involving search query generation and paper search using those queries left P10 with “more faith that I was getting a good coverage of the field.” P3 was impressed by Cocoa’s performance on thematic summarization steps: “[Cocoa] did a really good job at grouping papers that went together.” P11 also appreciated Cocoa’s speed at literature search and synthesis and found it conducive to early-stage explorations: “if I were to come up with a summary of 20 papers, it takes probably a day, but [in Cocoa] it’s around a minute which is really useful to formulate early-stage research ideas.” Beyond literature search, the structure of the plan was also useful to “break down tasks and provide a sense of progress” (P16) and “identifying important actionable steps I could take with a research problem” (P12). Overall, participants found Cocoa valuable in their work, with P3 in particular sharing that “I have an entire tab in my Google doc where we have stuff from Cocoa.”
7.2.4. Additional desiderata
Participants identified several desiderata for Cocoa after spending more time with it. Some participants wanted to clearly see what context Cocoa was used to generate plans and execute plan steps. For example, P10 was unsure whether the system was “using the context from my other plans to generate the [other] plan.” The system indeed was, but did not visualize relevant context in the document. P3, P15, and P16 all wanted more transparency into the shared context pool (Section 4.2) used across all steps and the specific context drawn upon by each step. Several participants desired tool interoperability, such that they could swap out Cocoa’s default tools for their own ones (e.g., an LLM finetuned on paper summaries instead of Cocoa’s paper summarizer). This way, interactive plans would serve as a “hub” for their custom tools. Finally, P5 took inspiration from Notion’s ability to nest pages to suggest nested plans, where plans can be organized and expanded in a hierarchical manner. These suggestions all provide valuable directions for future work.
8. Discussion